# The Repeat Offenders: Characterizing and Predicting Extremely Bug-Prone Source Methods

### v1

@@ 4.3. RQ3: Predicting the ExtremelyBuggy Methods
In this section, we investigate whether ExtremelyBuggy methods can be predicted at the time of their creation. Inspired by prior research (mo2022exploratory; pascarella_performance_2020; chowdhury_good_2025), we employ five machine learning models of increasing complexity: Logistic Regression (hosmer2013applied), Decision Tree (song2015decision), Random Forest (rigatti2017random), AdaBoost (schapire2013explaining), and a Multi-Layer Perceptron (MLP) model (murtagh1991multilayer). Our goal is to systematically evaluate the predictive power of both simple and complex models, assessing whether more sophisticated algorithms offer a significant advantage in identifying buggy code early in the development lifecycle. All models were implemented using the default parameters provided by the scikit-learn library. This choice was motivated by two factors: (i) hyperparameter tuning is computationally expensive under the leave-one-out validation setting—the evaluation approach used in this paper—as it requires training a separate model for each test instance, and (ii) prior similar studies have found that default settings often yield competitive performance (chowdhury_good_2025; pascarella_performance_2020). Moreover, using default parameters enhances the reproducibility of our study and facilitates future replication efforts.
@@
To build and train these models, we first merged the Buggy and NotBuggy methods into a single category: not ExtremelyBuggy. This transformation enables a binary classification task, allowing the models to specifically focus on distinguishing ExtremelyBuggy methods from all others. The models were trained using the 14 code metrics computed for each method at the time of its introducing commit. Making predictions by relying solely on metrics available immediately after a method is created allows us to assess the feasibility of detecting ExtremelyBuggy methods as early as possible in the software development process.
@@
Due to the highly imbalanced nature of our dataset, we do not use metrics such as accuracy to evaluate the performance of the machine learning models, as they could overstate the performance of our models. Instead, we use precision and recall, F-measure (F1), Area Under the Curve (AUC) score, and the Mathews Correlation Coefficient (MCC). AUC ranges from 0 to 1, with a score of 0.5 being no better than random classification and 1 being perfect. MCC ranges from -1 to 1, with 0 being about random classification and a score of 1 being perfect prediction.
@@
We assess prediction performance using a leave-one-out approach at the project level. In each iteration, all methods from one project are used as the test set, while methods from all other projects serve as the training data. This process is repeated for every project, ensuring that each serves as the test set exactly once. This evaluation strategy is motivated by two key factors: (i) it prevents data leakage between training and test sets (pascarella_performance_2020; chowdhury_method-level_2024), and (ii) it allows us to assess how well the models generalize across different projects. To address the data imbalance problem—specifically, the significantly smaller number of ExtremelyBuggy methods—we apply both undersampling and oversampling techniques using modules provided by the scikit-learn library. While the overall prediction performance remains similar across these strategies, undersampling yields slightly better results.
Figure 5 presents various metrics for two selected algorithms using the undersampling approach. The results indicate that recall is substantially higher than precision, suggesting that while the machine learning models are effective at identifying most ExtremelyBuggy methods, they tend to generate a high number of false positives. Most project-level evaluations fall within a similar range, but a few notable outliers stand out. For example, at least one project achieved perfect precision (1.00), despite approximately 95% of projects having precision below 0.50. Likewise, although around 90% of projects achieved recall scores above 0.5, there was at least one project with a recall below 0.25.
@@
Figure 5. Results for two of the select machine learning algorithms for the High Recall dataset. Results are very similar with other algorithms, and the results from the other algorithms can be viewed in our publicly shared repository  [^11] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/AdditionalFigures/LeaveOneOutTraining. To enhance readability, we use less number of markers than the original number of projects.:
RQ3 Summary: Predicting ExtremelyBuggy methods at inception using simple code metrics and standard machine learning algorithms remains highly challenging. While prediction performance varies across projects, the overall results are not yet reliable enough for practical use.
@@ 4.4. RQ4: Observable Characteristics of ExtremelyBuggy Methods
The futility of accurate prediction led us to manually characterize the ExtremelyBuggy methods to provide insights for both practitioners and the research community. We employed the thematic analysis (fereday2006demonstrating) approach to find themes of the ExtremelyBuggy methods. Thematic analysis involves identifying general patterns (or ”themes”) within the data. These themes are created by labeling the data with codes that capture and characterize individual instances of the data and later identifying themes within the codes. Our thematic analysis approach followed the guidelines of Cruzes et. al (CruzesD.S.2011RSfT), which provides a set of steps for thematic analysis in software engineering research. The approach is described as follows.
Data extraction.For this analysis, we’ve chosen the ExtremelyBuggy methods from only the HighPrecision dataset. The reasons behind this are: (i) it contains only 287 unique methods, allowing for thorough manual analysis, and (ii) it has the potential of minimal false positives compared to the other datasets. However, during our review, we still encountered some falsely labeled bugs—mainly due to the inclusion of the keyword flaw in commit messages (e.g., “fixing quality flaws”), which sometimes referred to non-functional changes. As a result, we excluded 22 such methods, leaving 265 ExtremelyBuggy methods and their evolution history for analysis.
Generating initial codes from the methods. A code is a label that represents some kind of specific, meaningful concept about the data. To help remove internal bias caused by preconceived ideas about the data, we used an open coding (strauss1998basics) approach to generate codes. This approach allowed us not to start with any code and create code intuitively, only creating new code when no existing codes fit the method.
To ensure consistency and correctness in the coding process, the first two authors independently coded 10 selected methods analyzing their history. Following this, the third and fourth authors—who have 2 and 10 years of industry experience, respectively—facilitated a discussion with the first two authors to review the initial codes. We realized that when manually analyzing a data sample, there are multiple dimensions one can take to characterize and assign ”codes” to it, as was also discussed in other studies (e.g., (jahan2025taxonomy)). This makes it difficult to generate very similar codes by two independent annotators. For example, in our case, one may choose to look at how a method is written, noting it is very large and hard to read, while another may look at the context it is written in, noting the method is used to parse a piece of text. As such, after the discussion session, we agreed to perform our manual analysis in accordance with the following three distinct dimensions:
@@
    Context codes. These codes represent the context a method is used. When generating these codes, we asked: What is the main purpose of this method (e.g., data handling)?
@@
    Bug codes. These codes are generated to represent the reasons for bugs in a method. When generating these codes, we asked for each bug in a method: Why did this bug occur (e.g., missing conditionals)?
The first two authors independently coded an additional set of 10 shared methods and produced highly similar codes. With growing confidence in their consistency, the remaining methods were divided between them for individual coding. While it is difficult to calculate agreement scores in open coding (cascio2019team), the authors conducted a validity check after analyzing all 265 methods. To assess consistency, each author randomly reviewed 30 methods originally coded by the other, and high agreement was observed—differences appeared in only 12 out of 60 cases. Importantly, these were not cases of conflicting codes; rather, one author had identified additional codes that the other had missed. For instance, one author noted the presence of self-admitted technical debt (SATD) (Potdar:2014; chowdhury_evidence_2025), which the other did not. We do not consider these additional codes a limitation. As discussed later, all identified codes, including the extra ones, were collaboratively reviewed by the two authors and synthesized into broader themes, ensuring the robustness and comprehensiveness of the analysis.
@@ 4.4.1. Results
In figure 6, we present an example of the complete thematic analysis process for one of our higher-order themes in the bug codes category and the percentages of methods that contain each of the codes and themes. The initial codes that the methods were labeled with are in the right column. In the middle column, we show the initial, more general themes that were identified after looking at the codes. For example, the codes logging and avoiding too many logs are combined into the more general theme of logging practices. Finally, the left column displays the higher-order theme, exception/error handling and logging that was identified using the two themes exception and error handling and logging practices. We combined logging and exception handling into a single broader category because both are commonly used to detect, respond to, and monitor erroneous behavior in program code. Together, they serve as essential tools for maintaining robustness, facilitating debugging, and ensuring system observability during failure scenarios (nakshatri2016analysis).
Figure 6. The taxonomy of one of the themes we identified in the bug themes category during our manual analysis of the ExtremelyBuggy methods. The right column shows initial codes, the middle shows more general themes, and the left contains the higher-order theme. Percentages are the number of ExtremelyBuggy methods containing that code/theme divided by the total number of ExtremelyBuggy methods. The complete taxonomies for all three themes (Bug, Context, and Visual) can be found in the shared repository.:
@@ 4.4.2. Visual Themes
  • Confusing Control Flow. This theme refers to methods where it is hard to follow the control flow of the method, for example if the method has many if/else statements or deeply nested code. An example is the processScope() method (105.json) from the intellij-community repository that has many many if/else statements and deeply nested code resulting in complicated bugs related to its control flow.
  • Abnormal Size. These methods are visually very large. This reinforces the common belief that very large methods are substantially more susceptible to bugs than smaller ones. The processEntry() method (35.json) from the jgit project is an example of a method that is abnormally large.
@@
  • SATD. Self-admitted technical debt (SATD) (potdar2014exploratory) occurs when developers acknowledge that there is future work that needs to be done on a method (e.g., with a TODO comment). An example is the parseTransition() method (21.json) from the argouml repository, where unresolved technical debts directly caused bugs in the future.
@@ 4.4.4. Bug Themes
In this dimension of themes, we focus on what was done to fix a bug, not what the method was doing (context) or how it looks (visual). Table 6 contains the final 10 themes we identified to explain types of bugs in ExtremelyBuggy methods.
@@
  • Variable Misuse and Naming. For these methods, bug-fixes were related to incorrect use and naming of variables. The method lookupValues() (44.json) from the pentaho-kettle repository provides an example where the wrong variable was passed to a method.
@@
  • Syntax, Type and Keyword Issues. These bugs occur due to typing issues, usually only requiring one keyword or type to be changed to be fixed. We also noticed that variables were often cast to the incorrect type, which caused the wrong methods to be called or being unable to be called. An example is the method supresses() (242.json) from the pmd repository.
@@
RQ4 Summary: ExtremelyBuggy methods often exhibit distinct visual characteristics (e.g., code smells such as large size or poor readability) and have frequent contextual roles (e.g., handling core logic). The underlying causes of bugs—such as missing edge cases or incorrect exception handling—are also commonly shared among these methods. These observations can help practitioners identify areas that warrant greater attention and support researchers in designing more effective features for future machine learning–based bug prediction models.
@@ 5. Discussion
Due to the importance of method-level bug prediction, significant research has been done in recent years (mo2022exploratory; giger_method-level_2012; chowdhury_method-level_2024; Shippey; ferenc_automatically_2020; pascarella_performance_2020). However, these earlier studies treated all bug-prone methods as the same, regardless of how many times a method was involved in bug fixes. In this study, we choose to explore the ExtremelyBuggy methods as understanding and capturing these methods would enable more optimized resource allocation. We found that ExtremelyBuggy methods are very small in numbers but can often account for the majority of bugs in a project (RQ1). This is encouraging because early optimization of these small numbers of methods can significantly reduce future maintenance burdens. Our analysis showed that ExtremelyBuggy methods are significantly larger, more complex, and less readable than both Buggy and NotBuggy methods (RQ2). However, ExtremelyBuggy methods proved very hard to predict at their inception, as machine learning models showed poor prediction performance (RQ3). Upon manual analysis, we found that the ExtremelyBuggy methods have common visual characteristics, contextual purposes, and similar fix types (RQ4).
@@ 5.1. Implications for Researchers
Many of the common themes we observed are not detectable from code metrics alone—e.g., HTTP and database operations, SATD, and abstract code operations. This implies that incorporating code embeddings as additional features might help improve the accuracy of these models. We also noticed that prediction performance is not uniform across all projects, as some projects exhibited unusually high or low accuracy. Future work should focus on optimal project selection for training models based on the project under test. Previous work in similar studies has shown the promise of optimal project selection during model training (chowdhury_method-level_2024; sun2021cfps).

### v2

@@ 4.3. RQ3: Predicting the ExtremelyBuggy Methods
In this section, we investigate whether ExtremelyBuggy methods can be predicted at the time of their creation. Inspired by prior research (Mo et al., 2022; Pascarella et al., 2020; Chowdhury, 2025), we employ five machine learning models of increasing complexity: Logistic Regression (Hosmer Jr et al., 2013), Decision Tree (Song and Lu, 2015), Random Forest (Rigatti, 2017), AdaBoost (Schapire, 2013), and a Multi-Layer Perceptron (MLP) model (Murtagh, 1991). Our goal is to systematically evaluate the predictive power of both simple and complex models, assessing whether more sophisticated algorithms offer a significant advantage in identifying ExtremelyBuggy methods early in the development lifecycle. All models were implemented using the default parameters provided by the scikit-learn library. This choice was motivated by three factors: (i) Parameter and hyperparameter tuning are computationally expensive under the leave-one-out validation setting—the evaluation approach used in this paper—as it requires training a separate model for each test instance; (ii) We also evaluated the models by randomly changing the parameter and hyperparameter values, and none of which improved the models’ accuracy. This also discouraged us from using the more expensive grid search approach. (iii) Prior similar studies found default settings yielded competitive accuracy (Chowdhury, 2025; Pascarella et al., 2020). Moreover, using default parameters enhances the reproducibility of our study and eases future replication efforts.
@@
To build and train these models, we first merged the Buggy and NotBuggy methods into a single category: not ExtremelyBuggy. This transformation enables a binary classification task, allowing the models to specifically focus on distinguishing ExtremelyBuggy methods from all others. The models were trained using the 14 code metrics computed for each method at the time of its introduction commit. Making predictions by relying solely on metrics available immediately after a method is created allows us to assess the feasibility of detecting ExtremelyBuggy methods as early as possible in the software development process.
@@
Due to the highly imbalanced nature of our dataset, we do not use metrics such as accuracy to evaluate the performance of the machine learning models, as they could overstate the performance of our models. Instead, we use precision, recall, F-measure (F1), Area Under the Curve (AUC) score, and the Mathews Correlation Coefficient (MCC). AUC ranges from 0 to 1, with a score of 0.5 being no better than random classification and 1 being perfect. MCC ranges from -1 to 1, with 0 being about random classification and a score of 1 being perfect prediction.
@@
We assess prediction performance using a leave-one-out approach at the project level. In each iteration, all methods from one project are used as the test set, while methods from all other projects serve as the training data. This process is repeated for every project, ensuring that each serves as the test set exactly once. This evaluation strategy is motivated by two key factors: (i) it prevents data leakage between training and test sets (Pascarella et al., 2020; Chowdhury et al., 2024), and (ii) it allows us to assess how well the models generalize across different projects. To address the data imbalance problem—specifically, the significantly smaller number of ExtremelyBuggy methods—we apply both undersampling and oversampling techniques using modules provided by the scikit-learn library. While the overall prediction performance remains similar across these strategies, undersampling yields slightly better results.
@@
Figure 6. Results for two of the select machine learning algorithms for the Balanced dataset. Results are very similar to those of other datasets and algorithms. Results from the other algorithms can be viewed in our publicly shared repository. [^10] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/AdditionalFigures/LeaveOneOutTraining
Results for two of the select machine learning algorithms for the High Recall dataset.
Figure 6 presents various metrics for two selected algorithms using the undersampling approach. The results indicate that recall is substantially higher than precision, suggesting that while the machine learning models are effective at identifying most ExtremelyBuggy methods, they tend to generate a very high number of false positives. Most project-level evaluations fall within a similar range, but a few notable outliers stand out. For example, at least one project achieved precision above 0.50, despite approximately 75% of projects having precision below 0.25. Likewise, although around 90% of projects achieved recall scores above 0.5, there was at least one project with a recall below 0.25.
RQ3 Summary Predicting ExtremelyBuggy methods at inception using simple code metrics and standard machine learning algorithms remains highly challenging. While prediction performance varies across projects, the overall results are not yet reliable enough for practical use.
@@ 4.4. RQ4: Observable Characteristics of ExtremelyBuggy Methods
The futility of accurate prediction led us to manually characterize the ExtremelyBuggy methods to provide insights for both practitioners and the research community. We employed the thematic analysis (Fereday and Muir-Cochrane, 2006) approach to find themes of the ExtremelyBuggy methods. Thematic analysis involves identifying general patterns (or ”themes”) within the data. These themes are created by labeling the data with codes that capture and characterize individual instances of the data and later identifying themes within the codes. Our thematic analysis approach followed the guidelines of Cruzes et. al (Cruzes and Dyba, 2011), which provides a set of steps for thematic analysis in software engineering research. The approach is described as follows.
Data extraction. For this analysis, we selected the ExtremelyBuggy methods from only the HighPrecision dataset. The reasons behind this are: (i) it contains only 287 unique methods, allowing for thorough manual analysis, and (ii) it has the potential of minimal false positives compared to the other datasets. However, during our review, we still encountered some falsely labeled bugs—mainly due to the inclusion of the keyword flaw in commit messages (e.g., “fixing quality flaws”), which sometimes referred to non-functional changes. As a result, we excluded 22 such methods, leaving 265 ExtremelyBuggy methods and their evolution history for analysis.
Generating initial codes from the methods. A code is a label that represents a specific, meaningful concept within the data. To minimize internal bias stemming from preconceived assumptions, we employed an open coding approach (Strauss and Corbin, 1998) to generate codes. Using this approach, we did not begin with predefined codes; instead, we derived codes directly from the data and created new codes whenever the already created codes did not adequately capture the content.
To ensure consistency and correctness in the coding process, the first two authors independently coded 10 selected methods. Following this, the third and fourth authors—who have 2 and 10 years of industry experience, respectively—facilitated a discussion with the first two authors to review the initial codes. We realized that when manually analyzing a data sample, there are multiple dimensions one can take to characterize and assign ”codes” to it, as was also discussed in other studies (e.g., (Jahan et al., 2025)). This makes it difficult to generate very similar codes by two independent annotators. For example, in our case, one may choose to look at how a method is written, noting it is very large and hard to read, while another may look at the context it is written in, noting the method is used to parse a piece of text. As such, after the discussion session, we agreed to perform our manual analysis in accordance with the following three distinct dimensions:
@@
    Context codes. These codes represent the context in which a method is used. When generating these codes, we asked: What is the main purpose of this method (e.g., data handling)?
@@
    Bug codes. These codes are generated to represent the reasons for bugs in a method. When generating these codes, we checked the bug fix and asked: Why did this bug occur (e.g., missing conditionals)?
The first two authors independently coded an additional set of 10 shared methods and produced highly similar codes. After more verification with the third and fourth authors and growing confidence in labeling consistency, the remaining methods were divided between the first and second authors for individual coding. While it is difficult to directly calculate agreement scores in open coding (cascio2019team), as two authors can generate similar codes with different names, the authors conducted a validity check after analyzing all 265 methods. To assess consistency, each author randomly reviewed 30 methods originally coded by the other, and high agreement was observed—differences appeared in only 12 out of 60 cases. Importantly, these were not cases of conflicting codes; rather, one author had identified additional codes that the other had missed. For instance, one author noted the presence of self-admitted technical debt (SATD) (Potdar:2014; Chowdhury et al., 2025), which the other did not. As discussed later, all identified codes, including the extra ones, were collaboratively reviewed by the two authors and synthesized into broader themes, ensuring the robustness and comprehensiveness of the analysis.
@@ 4.4.1. Results
In figure 7, we present an example of the complete thematic analysis process for one of our higher-order themes in the bug codes category and the percentages of methods that contain each of the codes and themes. The initial codes that the methods were labeled with are in the right column. In the middle column, we show the initial, more general themes that were identified after looking at the codes. For example, the codes logging and avoiding too many logs are combined into the more general theme of logging practices. Finally, the left column displays the higher-order theme, exception/error handling and logging, that was identified using the two themes exception and error handling and logging practices. We combined logging and exception handling into a single broader category because both are commonly used to detect, respond to, and monitor erroneous behavior in program code. Together, they serve as essential tools for maintaining robustness, facilitating debugging, and ensuring system observability during failure scenarios (nakshatri2016analysis).
Figure 7. The taxonomy of one of the themes we identified in the bug themes category during our manual analysis of the ExtremelyBuggy methods. The right column shows initial codes, the middle shows more general themes, and the left contains the higher-order theme. Percentages are the number of ExtremelyBuggy methods containing that code/theme divided by the total number of ExtremelyBuggy methods. The complete taxonomies for all three themes (Bug, Context, and Visual) can be found in the shared repository. [^12] https://github.com/SQMLab/ExtremelyBuggyPublicData/tree/main/AdditionalFigures/ThematicAnalysisTaxonomy
The taxonomy of one of the themes we identified in the bug themes category during our manual analysis
@@ 4.4.2. Visual Themes
  • Confusing Control Flow. This theme refers to methods where it is hard to follow the control flow of the method, for example, if the method has many if/else statements or deeply nested code. An example is the processScope() method (105.json) from the intellij-community repository that has many if/else statements and deeply nested code, resulting in complicated bugs related to its control flow.
  • Abnormal Size. These methods are visually very large. This reinforces the common belief that very large methods are substantially more susceptible to bugs than smaller ones (Gil and Lalouche, 2017; Chowdhury et al., 2024, 2022a). The processEntry() method (35.json) from the jgit project is an example of a method that is abnormally large.
@@
  • SATD. Self-admitted technical debt (SATD) (Potdar and Shihab, 2014) occurs when developers acknowledge that there is future work that needs to be done on a method (e.g., with a TODO comment). An example is the parseTransition() method (21.json) from the argouml repository, where unresolved technical debts directly caused bugs in the future.
@@ 4.4.4. Bug Themes
In this dimension of themes, we focus on what was done to fix a bug, not what the method is doing (context) or how it looks (visual). Table 6 contains the final 10 themes we identified to explain types of bugs in ExtremelyBuggy methods.
@@
  • Variable Misuse and Naming. For these methods, bug-fixes were related to the incorrect use and naming of variables. The method lookupValues() (44.json) from the pentaho-kettle repository provides an example where the wrong variable was passed to a method.
@@
  • Syntax, Type and Keyword Issues. These bugs occur due to typing issues, usually only requiring one keyword or type to be changed to be fixed. We also noticed that variables were often cast to the incorrect type, which caused the wrong methods to be called or were unable to be called. An example is the method supresses() (242.json) from the pmd repository.
@@
RQ4 Summary ExtremelyBuggy methods often exhibit distinct visual characteristics (e.g., code smells such as large size or poor readability) and have frequent contextual roles (e.g., handling core logic). The underlying causes of bugs—such as missing edge cases or incorrect exception handling—are also commonly shared among these methods. These observations can help practitioners identify areas that warrant greater attention and support researchers in designing more effective features for future machine learning–based bug prediction models.
@@ 5. Discussion
Due to the importance of method-level bug prediction, significant research has focused on this topic in recent years (Mo et al., 2022; Giger et al., 2012; Chowdhury et al., 2024; Shippey et al., 2016; Ferenc et al., 2020; Pascarella et al., 2020). However, these earlier studies treated all bug-prone methods as the same, regardless of how many times a method was involved in bug fixes. In this study, we choose to explore the ExtremelyBuggy methods as understanding and capturing these methods would enable more optimized resource allocation. We found that ExtremelyBuggy methods are very small in numbers but can often account for a large number of bugs in a project (RQ1). This is encouraging because early optimization of these small numbers of methods can significantly reduce future maintenance burdens. Our analysis showed that ExtremelyBuggy methods are significantly larger, more complex, and less readable than both Buggy and NotBuggy methods (RQ2). Despite these differences, ExtremelyBuggy methods are very challenging to predict at their inception, as machine learning models showed poor prediction performance (RQ3). However, with manual analysis, we found that the ExtremelyBuggy methods have common visual characteristics, contextual purposes, and similar root causes of bugs (RQ4).
@@ 5.1. Implications for Researchers
Many of the common themes we observed are not detectable from code metrics alone—e.g., HTTP and database operations, SATD, and abstract code operations. This implies that incorporating code embeddings as additional features might help improve the accuracy of these models—this, however, will be an expensive process given the enormous amount of methods we have. We also noticed that prediction performance is not uniform across all projects, as some projects exhibited unusually high or low accuracy. This implies that future research can focus on optimal project set selection for training models based on the project under test. Previous work in similar studies has shown the promise of optimal project selection during model training (Chowdhury et al., 2024; Sun et al., 2021).
One of the major challenges in method-level bug prediction (Chowdhury et al., 2024), and thus in predicting ExtremelyBuggy methods, is the presence of tangled commits. Although a significant amount of research (muylaert_research_2018; partachi_flexeme_2020; li_utango_2022; xu_detecting_2025) has focused on untangling tangled code changes, none of these approaches addresses the problem at the method-level granularity. Only recently, Opu et al. (Opu et al., 2026) evaluated the effectiveness of Large Language Models for detecting tangling in method-level bug-fix commits and demonstrated significant success. Future research should investigate whether applying untangling as a preprocessing step can further improve the accuracy reported in our study.
Another challenge we encountered was the data imbalance issue, as there were very few ExtremelyBuggy samples compared to other classes. This is a well-known problem that makes robust model training particularly challenging (Manahel:2025). Although we attempted to mitigate this issue using simple oversampling and undersampling techniques, future research could explore more advanced RAG-based augmentation approaches (daneshvar2025vulscriber) to improve minority-class representation and overall model robustness.